OpenStudy (anonymous):

How do you graph f(x)=x(x-1)?

OpenStudy (sohailiftikhar):

just put the values of x in function and get the value of f(x) which is equal to y first put x=0 you will get y=0 (0,0) x=1 you will get y=0 (1,0) now put x=2 then y=2 (2,2) put x=3 then y=6 (3,6) etc

OpenStudy (campbell_st):

its a parabola with zeros at 0, 1 and a line of symmetry of x =1/2 its concave up so you can use a table of values x: -3 -2 -1 0 1 2 3 y : 12 6 2 0 0 2 6

OpenStudy (sohailiftikhar):

now mark these points on x and y axes